[Hallc_running] [New Logentry] RC update -July 30, 2025
arunin at jlab.org
arunin at jlab.org
Wed Jul 30 12:15:02 EDT 2025
Logentry Text:
--
Data taking proceeded smoothly until the Owl shift, when the HMS magnets (Q1, Q3, and Dipole) interlocked due to a trip of the Hall B solenoid (https://logbooks.jlab.org/entry/4430439, https://logbooks.jlab.org/entry/4430352). We also encountered several target-related alarms.
Chandan reported issues with the cryoflow and heater power of the Loop2-LH2 target.
The SHMS dipole remains in manual mode due to a communication issue. The associated alarm has been silenced.
We are currently collecting data at the z = 0.5 π+ SIDIS setting. Beam studies for Hall C were completed earlier this morning.
Today, Wednesday 7/30, is a beam studies/maintenance day. The hall will enter controlled access to allow Ellen to work on the SHMS dipole magnet.
We will hold our regular RC meeting today at 2:30 PM. The Zoom link is provided below.
https://jlab-org.zoomgov.com/j/1618227115?pwd=8JkhFg6YNesVc3HNgWjwUSIKIu...
Meeting ID: 161 822 7115
Passcode: 253463
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4430726
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250730/078f558a/attachment.htm>
More information about the Hallc_running
mailing list